Dillon Products manufactures various machined parts to customer specifications. The company uses a job-order costing system and applies overhead cost to jobs on the basis of machine-hours. At the beginning of the year, the company used a cost formula to estimate that it would incur $4,262,400 in manufacturing overhead cost at an activity level of 576,000 machine-hours.

The company spent the entire month of January working on a large order for 12,100 custom-made machined parts. The company had no work in process at the beginning of January. Cost data relating to January follow:

  1. Raw materials purchased on account, $322,000.
  2. Raw materials used in production, $266,000 (80% direct materials and 20% indirect materials).
  3. Labor cost accrued in the factory, $150,000 (one-third direct labor and two-thirds indirect labor).
  4. Depreciation recorded on factory equipment, $63,100.
  5. Other manufacturing overhead costs incurred on account, $85,300.
  6. Manufacturing overhead cost was applied to production on the basis of 40,620 machine-hours actually worked during the month.
  7. The completed job for 12,100 custom-made machined parts was moved into the finished goods warehouse on January 31 to await delivery to the customer. (In computing the dollar amount for this entry, remember that the cost of a completed job consists of direct materials, direct labor, and applied overhead.)

Required:

1. Prepare journal entries to record items (a) through (f) above [ignore item (g) for the moment].

2. Prepare T-accounts for Manufacturing Overhead and Work in Process. Post the relevant items from your journal entries to these T-accounts.

3. Prepare a journal entry for item (g) above.

4. If 10,600 of the custom-made machined parts are shipped to the customer in February, how much of this job’s cost will be included in cost of goods sold for February?
Answer #1

Answers:

1.

(a).

General Journal Debit Credit
Raw Material Inventory       322,000
Accounts Payable 322,000

(b).

Work in Process       212,800
Manufacturing Overhead         53,200
Raw Material Inventory              266,000

(c).

Work in Process         50,000
Manufacturing Overhead       100,000
Wages Payable              150,000

(d).

Manufacturing Overhead         63,100
Accumulated Depreciation                63,100

(e).

Manufacturing Overhead         85,300
Accounts Payable                85,300

(f).

Work in Process       300,588
Manufacturing Overhead              300,588

Calculation

Predetermined overhead rate = 4,262,400 / 576,000 = $7.40 per machine hour.

So, 7.40 x 40,620 = 300,588

2.

Manufacturing Overhead
Raw Material Inventory 53,200 Work in Process 300,588
Wages Payable 100,000 Balance under-applied) 1,012
Accumulated Depreciation 63,100
Accounts Payable 85,300
301,600 301,600
Work in Process
Raw Material Inventory 212,800 Finished Goods 563,388
Wages Payable 50,000
Manufacturing Overhead 300,588
563,388 563,388

3.

(g).

Finished Goods 563,388
Work in Process    563,388

Calculation

Total costs of job = 212,800 + 50,000 + 300,588 = 563,388

4.

Unit Product Cost = 563,388 / 12,100 = $46.56 per unit

So, the portion of this job’s costs that would be included in February’s cost of goods sold is:

= 10,600 units × $46.56 per unit = $493,546.51